tina has 70 coins,all quarters and dimes. there are 30 more quarters than dimes. find the number of each type of coin.

Answer by checkley79(3341)

Q+D=70
Q=D+30
(D+30)+D=70
2D+30=70
2D=70-30
2D=40
D=40/2
D=20 ANS. FOR THE NUMBER OF DIMES.
Q+20=70
Q=70-20
Q=50 ANS. FOR THE NUMBER OF QUARTERS.
PROOF:
20=50=70
70=70